OpenStudy (anonymous):

find the surface area of a cone with the raduis 8.8 cm and slant height of 7.1 cm , to the nearest square unit , whats the answer i wanted to check my work 636cm^2 ?

OpenStudy (anonymous):

or 683cm^2

OpenStudy (anonymous):

No. It's wrong. SA = πr^2 + πrl SA = π(8.8)^2 + π(8.8)(7.1) SA = 77.44π + 62.48π SA = 139.92π SA = 439.3488 cm^2

OpenStudy (anonymous):

ughh , mann

OpenStudy (anonymous):

This the same problem as before right?

OpenStudy (anonymous):

You probably just multiplied incorrectly or something.

OpenStudy (anonymous):

yea , i was checking my work

OpenStudy (anonymous):

but thanks a lot (:

OpenStudy (anonymous):

You're welcome :)